OpenStudy (anonymous):

Can someone tell me if I'm right: Simplify the given expression to rational exponent form, justify each step by identifying the properties of rational exponents used. 1 over the cube root of the quantity of x to the negative sixth power = 3squrt x^-6= x^-6/3

OpenStudy (anonymous):

\[1/3\sqrt{-6}=3\sqrt{-6}=x^{-6}/3\]

OpenStudy (anonymous):

that's suppose to be x^-6 in the squrt

OpenStudy (e.mccormick):

\(\dfrac{1}{\sqrt[3]{x^{-6}}}\) You can't just take it out from under the 1.

OpenStudy (anonymous):

okay...so how so I solve what the question is asking?

OpenStudy (e.mccormick):

Deal with the cube first, then it will be in a form where you can use the fact that \(\frac{1}{x^{-1}}=x\) to get it out from the bottom of the fraction.

OpenStudy (anonymous):

I have to make it a rational exponent though...im kind of confused

OpenStudy (e.mccormick):

That is a rational exponent, when you deal with the cube root. Know what a root is as an exponent? \(\sqrt{x}=x^{\frac{1}{2}}\) So you do the same thing with the cube root, but remembering that there alreadu is the -6 power on x.

OpenStudy (anonymous):

\[ x \frac{ 1 }{ -6}?\]

OpenStudy (e.mccormick):

Let me show it with some variables. \(\LARGE \sqrt[b]{x^a}=x^{\frac{a}{b}}\)

OpenStudy (e.mccormick):

See how the b is on the bottom of the fraction in the rational exponent?

OpenStudy (anonymous):

\[x \frac{ -6}{ 3 }\]

OpenStudy (anonymous):

Like that?

OpenStudy (e.mccormick):

Yah. The - can be out front, but on top is OK too.

OpenStudy (e.mccormick):

Now you can use the \(\frac{1}{x^{-1}}=x^1\) rule.

OpenStudy (anonymous):

okay you lost me

OpenStudy (anonymous):

\[x1-\frac{ 6 }{ 3 }\]

OpenStudy (e.mccormick):

A negatibve exponent causes something to invert . Some examples right on the top of this: http://www.purplemath.com/modules/simpexpo2.htm

OpenStudy (anonymous):

so x will have a negative exponent?

OpenStudy (e.mccormick):

No, you want to get rid of the negative part of the exponent. That makes it go from one side of the fraction to the other. In this case, it gets rid of the 1 over part because there is nothing else to the fraction.

OpenStudy (anonymous):

okay so it will still be: \[x \frac{ 6 }{ 3}\] Just know negative sign

OpenStudy (e.mccormick):

Correct. No negative sign and no one over. And finally, you can simplify the rational exponent. They work like regular fractions.

OpenStudy (e.mccormick):

Then it will be in fully simplified form.

OpenStudy (anonymous):

okay...um so do I do 6 divided by 3...sorry i'm really bad at math

OpenStudy (e.mccormick):

Yes. 6 divided by 3.

OpenStudy (anonymous):

so x2

OpenStudy (e.mccormick):

\(x^2\) Yep. That is the final form.

OpenStudy (anonymous):

Yeah! Thank you so much for explaining it to me! ^-^

OpenStudy (e.mccormick):

When you type up your explanation, use x^(-6/3) because the ( ) make it clear what you mean.

OpenStudy (anonymous):

Okay I'll do that Thanks!

OpenStudy (e.mccormick):

np. have fun!